OpenStudy (anonymous):

solve equation by square root property (m + 12) sq root 2 = -9

OpenStudy (anonymous):

@geerky42 @jim_thompson5910

OpenStudy (tkhunny):

Please describe the "Square Root Property".

OpenStudy (anonymous):

@geerky42 it's similar to the last one

OpenStudy (anonymous):

is it 9-3 or 9 + 3 ?

geerky42 (geerky42):

Do the same thing I showed you. Take square root of both sides. Note that your answer will contain imaginary number.

OpenStudy (anonymous):

@geerky42 but wait whats the sq root of 12?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (tkhunny):

Are you going to keep looking for volunteers until someone does the work for you? Show YOUR work and demonstrate some competence in the material.

OpenStudy (anonymous):

excuse me? I'm asking for help, get out of here I didn't ask you

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (tkhunny):

It would help if you would answer my question.

OpenStudy (anonymous):

why? I simply did not ask you

OpenStudy (anonymous):

I know the answer, but I'm not going to answer now with the way you just expressed yourself to me..sorry

OpenStudy (tkhunny):

The help you need may not be apparent to you. Demonstrate your answer. I'm not being mean or rude, but I understand this is a common misconception with some difficult honesty. Show your work. Seriously, it will help you.

OpenStudy (anonymous):

I did! lol

OpenStudy (anonymous):

is it 9-3 or 9 + 3 ?

OpenStudy (anonymous):

that's how far i've gotten :)

OpenStudy (tkhunny):

That's not an answer. Neither of those makes any sense. How did you get those? What was your plan? What, exactly, is the "Square Root Property"? Are you SURE you have written to problem correctly? So far, it looks like this: \((m + 12)\sqrt{2} = -9\) Is that the right problem? As written, we do NOT need the Square Root Property. I'm a little puzzled.

OpenStudy (anonymous):

? That's not the problem..the 2 is up high, it dosen't have the symbol

OpenStudy (anonymous):

i'm not SAYING that's the answer, i'm saying that's how far i got, sq rooting the numbers :P

OpenStudy (tkhunny):

That's part of what we can learn when we see your work. If you mean \(({m+12})^{2} = -9\), you can write it so that it can be understood more readily. Maybe (m+12)^2 = -9. Does that make sense?

OpenStudy (anonymous):

that's the right problem

OpenStudy (anonymous):

so wouldnt i do 12 x 12?

OpenStudy (anonymous):

or 12 x 2

OpenStudy (tkhunny):

If you had written "sq 2", I may have interpreted it correctly. You need to see that your use of the word "root" was confusing. The Square Root Property is this: If \(x^{2} = a\;then\;x = \pm\sqrt{a}\). By this property, \((m+12)^{2} = -9\;leads\;to\;m+12 = \pm\sqrt{-9}\) Do you see how this is a DIRECT application of the property? That is what we need to see. That is what you need to understand. Apply the property. See it. Understand what it does.

OpenStudy (anonymous):

okay? so

OpenStudy (tkhunny):

So, show your work. What's next? I would suggest subtraction?

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (tkhunny):

I would also suggest simplifying \(\pm\sqrt{-9}\).

OpenStudy (anonymous):

i know they're opposite so i do 12 - 9?

OpenStudy (anonymous):

i'm graduating next week from high school, I just have to finish this, then i'm done. I'm not studying this in college, so lol